OpenStudy (josh0404):

what is the largest function value of f(x) = -x^2+4?

jimthompson5910 (jim_thompson5910):

hint: look at the y coordinate of the vertex

OpenStudy (josh0404):

how is it sir, i dont know how to solve?

jimthompson5910 (jim_thompson5910):

do you know how to graph?

jimthompson5910 (jim_thompson5910):

I would use a graphing tool like desmos https://www.desmos.com/calculator

OpenStudy (josh0404):

yes sir

OpenStudy (josh0404):

how to solve this one sir, please help me?

jimthompson5910 (jim_thompson5910):

ok graph it and see where the peak point is

jimthompson5910 (jim_thompson5910):

type in `-x^2+4`

OpenStudy (josh0404):

the graph is on maximum pick sir

OpenStudy (josh0404):

because a is negative that is why the graph is downward

jimthompson5910 (jim_thompson5910):

what's the highest point you see?

OpenStudy (josh0404):

4 sir

jimthompson5910 (jim_thompson5910):

(0,4) is the point, but y = 4 is the answer, yes 4 is the largest output possible

OpenStudy (josh0404):

sir can you give me tips, how to analyze question easily?

jimthompson5910 (jim_thompson5910):

you can think of y = -x^2 + 4 as y = -1(x-0)^2 + 4 then compare that to vertex form y = a(x-h)^2 + k vertex = (h,k) = (0,4)

OpenStudy (danjs):

The general standard quadratic is the parabola shape. the equation of standard form is y = ax^2 + bx + c you were right with the first term, ax^2, and the sign determins if the parabola opens upwards or downwards in the y direction. A parabola like this will have either a max value , or a min value depending which way it opens. For this prob there is a negative -ax^2, and it opens downwards. So it goes downwards forever to infinity, and so no minimum value. The maximum value will be at the vertex of the parabola.

OpenStudy (mathstudent55):

You have \(f(x) = -x^2+4\), and you want to find the maximum value of f(x). Notice you have x^2. No matter what value you use for x, x^2 will always be non-negative. x^2 will have a minimum value of zero. x^2 is zero or greater. The expression has -x^2, so -x^2 will be non-positive. The maximum value of -x^2 is zero. -x^2 will be zero or lower. Since the maximum value of -x^2 is zero, the maximum value of -x^2 + 4 is 4.